### v1

# Theory of polarization-switchable electrical conductivity anisotropy in nonpolar semiconductors

Abstract
The anisotropic propagation of particles is a fundamental transport phenomenon in solid state physics. As for crystalline semiconductors, the anisotropic charge transport opens novel designing routes for electronic devices, where the electrical manipulation of anisotropic resistance provides essential guarantees. Motivated by the concept of anisotropic magnetoresistance, we develop an original theory on the electrically manipulatable anisotropic electroresistance. We show that when gaining electric polarization, nonpolar semiconductors can showcase anisotropic electrical conductivities between two perpendicular directions and the electrical conductivity anisotropy (ECA) is switchable by flipping the polarization. By symmetry analysis, we identify several point groups hosting the polarization-switchable ECA. These point groups simultaneously enable polarization-reversal induced conductivity change along specific directions, akin to the tunnelling electroresistance in ferroelectric tunnel junctions. First-principles-based conductivity calculations predict that AlP and NaSrP are two good materials having such exotic charge transport. Our theory can motivate the design of intriguing anisotropic electronic devices (e.g., anisotropic memristor and field effect transistor).
Introduction. The anisotropic transport in crystalline materials — for instance, the anisotropic propagation of magnons Cui et al. (2024); Qi et al. (2023), polaritons Luo et al. (2023), electrons or holes Wang et al. (2019a); Wadley et al. (2016); Jungwirth et al. (2016); Cording et al. (2024); Dai et al. (2022); Zeng et al. (2020); Wang et al. (2019b); Liu et al. (2015); Xia et al. (2014); Benítez et al. (2018) — not only is of fundamental interest to solid state physics, but offers extensive development strategies for electronic, optoelectronic and spintronic devices Zhao et al. (2020); Xia et al. (2014); Baltz et al. (2018); Ritzinger and Výborný (2023); Jungwirth et al. (2016); Benítez et al. (2018). Mostly, the design of such devices relies on the electrical or magnetic manipulation of anisotropic transport Qi et al. (2023); Luo et al. (2023); Wang et al. (2019a); Wadley et al. (2016); Jungwirth et al. (2016); Benítez et al. (2018). A profound example is the anisotropic magnetoresistance that occurs in magnets, where longitudinal electrical resistance therein depends on the direction of magnetic order Baltz et al. (2018); Ritzinger and Výborný (2023); Jungwirth et al. (2016); Walter et al. (2014). In ferromagnetic materials, magnetic field manipulates the magnetization direction, and changes the electrical resistance accordingly Baltz et al. (2018); Jungwirth et al. (2016). This gives birth to early-stage magnetic recording devices Baltz et al. (2018); Ritzinger and Výborný (2023), and yields magnetic sensors extensively deployed in automotive, aerospace, and biomedical realms Ritzinger and Výborný (2023).
The anisotropic magnetoresistance is reminiscent of the conceptual anisotropic electroresistance. Previous works reveal the occurrence of tunnelling electroresistance, a counterpart of tunnelling magnetoresistance, in ferroelectric tunnel junctions, with the electroresistance characterized by the electrical resistance change induced by the reversal of electric polarization Jia et al. (2024); Wu et al. (2020); Xi et al. (2017); Wen et al. (2013); Zhuravlev et al. (2005); Ding et al. (2021); Liu et al. (2016); Hernandez-Martin et al. (2020); Klyukin et al. (2018). With regard to anisotropic electroresistance, a particularly interesting research avenue is to discover crystalline semiconductors with polarization-switchable electrical conductivity anisotropy (ECA). Explicitly, it is desired to find nonpolar semiconductors exhibiting $P$-dependent anisotropic conductivities between two orthogonal $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions [i.e., $\sigma_{\chi,\chi}(P)\neq\sigma_{\chi^{\prime},\chi^{\prime}}(P)$], so that
$\sigma_{\chi,\chi}(P)-\sigma_{\chi^{\prime},\chi^{\prime}}(P)=-\sigma_{\chi,\chi}(-P)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P),$ (1)
where $P$ is an electric polarization induced by external electric field. Equation (1) guarantees that the ECA is manipulable and switchable by external electric field, which promises the design of anisotropic electronic devices (e.g., anisotropic memristor and field effect transistor) Zhao et al. (2020); Wang et al. (2019a); Liu et al. (2015). Besides that, Eq. (1) hints a possibility for a polarization-reversal induced conductivity change along $\bm{\chi}$ (or $\bm{\chi^{\prime}}$) direction in crystalline materials  [^1] According to Eq. (1), the $\sigma_{\chi,\chi}(P)-\sigma_{\chi,\chi}(-P)$ conductivity difference along $\bm{\chi}$ direction equals $\sigma_{\chi^{\prime},\chi^{\prime}}(P)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P)-2\sigma_{\chi,\chi}(-P)$, which is not necessarily zero. The similar logic applies to the conductivity difference along $\bm{\chi^{\prime}}$ direction. In next section, we shall discuss this point in details., mimicking the tunnelling electroresistance in ferroelectric tunnel junctions. In such sense, semiconductors with polarization-switchable ECA may serve as alternates for ferroelectric tunnel junctions in device design.
This work aims to provide guidelines for discovering the aforementioned elusive semiconductors. To this end, we develop a polarization-dependent Boltzmann transport theory within effective mass approximation. Then, we perform symmetry analysis for nonpolar crystallographic point groups, derive their symmetry-adapted energy dispersions, and extract the point groups that host polarization-switchable ECA. Simultaneously, our extracted point groups enable the polarization-reversal driven conductivity change. This is well supported by first-principles-based electrical conductivity calculations, which further predict semiconducting AlP and NaSrP as two representative materials with such exotic transport phenomena. Our theory emphasizes the potential of various nonpolar semiconductors (including the celebrated zinc-blende type semiconductors) in designing devices based on anisotropic charge transport.
Theory of polarization-switchable ECA. We start by developing the polarization-dependent Boltzmann transport theory. Under the approximation of constant relaxation time $\tau$, the longitudinal electrical conductivity (at zero Kelvin), contributed by a polarization-dependent $\epsilon(\mathbf{k},P)$ energy dispersion, is given by Tsirkin (2021); Scheidemantel et al. (2003); Železný et al. (2021); Zhao et al. (2024)
$\frac{\sigma_{\chi,\chi}(P,\mu)}{\tau}\propto\iiint\left[\frac{\partial\epsilon(\mathbf{k},P)}{\partial k_{\chi}}\right]^{2}\delta(\epsilon(\mathbf{k},P)-\mu)\,d^{3}\mathbf{k},$ (2)
where $\delta$, $\mu$, $P$, and $\mathbf{k}\equiv k_{x}\mathbf{x}+k_{y}\mathbf{y}+k_{z}\mathbf{z}$  [^2] We employ the $xyz$ Cartesian coordinate system with $x$, $y$, and $z$ axes being perpendicular to each other. As a convention, we use $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ to denote the unit vectors along $x$, $y$, and $z$ axes, respectively. denote Dirac’s $\delta$ function, chemical potential, electric polarization, and electronic momentum, respectively. Writing the electrical conductivity as $\sigma_{\chi,\chi}(P,\mu)$ emphasizes its dependence on electric polarization and chemical potential.
To demonstrate our basic ideas, we work with a simplified energy dispersion
$\epsilon(\mathbf{k},P)=\alpha_{P}k_{x}^{2}+\beta_{P}k_{y}^{2}+\gamma_{P}k_{z}^{2},$ (3)
with $\alpha_{P}$, $\beta_{P}$, and $\gamma_{P}$ depending on electric polarization. For simplicity too, we explore the ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions by assuming that $\alpha_{P}$, $\beta_{P}$, and $\gamma_{P}$ are all positive numbers. The evaluation of Eqs. (2) and (3) yields
$\frac{\sigma_{x,x}(P,\mu)}{\tau}\propto\frac{\alpha_{P}\mu^{\frac{3}{2}}}{\sqrt{\alpha_{P}\beta_{P}\gamma_{P}}},\frac{\sigma_{y,y}(P,\mu)}{\tau}\propto\frac{\beta_{P}\mu^{\frac{3}{2}}}{\sqrt{\alpha_{P}\beta_{P}\gamma_{P}}}.$ (4)
Up to first order in $P$, we expand $\alpha_{P}$, $\beta_{P}$, and $\gamma_{P}$ as $\alpha_{P}=\alpha_{0}+\alpha_{1}P$, $\beta_{P}=\beta_{0}+\beta_{1}P$, and $\gamma_{P}=\gamma_{0}+\gamma_{1}P$. This implies that
$\frac{\sigma_{x,x}(P,\mu)-\sigma_{y,y}(P,\mu)}{\tau}\propto\frac{(\alpha_{0}-\beta_{0})+(\alpha_{1}-\beta_{1})P}{\sqrt{\alpha_{P}\beta_{P}\gamma_{P}}}\mu^{\frac{3}{2}}.$ (5)
To achieve the polarization-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions, it is required that $\alpha_{0}=\beta_{0}$, $\alpha_{1}\neq\beta_{1}$  [^3] Our theory is established under the first-order approximation with respect to electric polarization $P$. Beyond such an approximation, this statement is not exactly true. For instance, $\alpha_{P}=\alpha_{0}+\alpha_{1}P+\alpha_{3}P^{3}$ and $\beta_{P}=\alpha_{0}+\beta_{1}P+\beta_{3}P^{3}$ may yield non-zero $\sigma_{x,x}(P,\mu)-\sigma_{y,y}(P,\mu)$, when $\alpha_{1}=\beta_{1}$ and $\alpha_{3}\neq\beta_{3}$., and $\alpha_{P}\beta_{P}\gamma_{P}=\alpha_{-P}\beta_{-P}\gamma_{-P}$ [see Eq. (1)]. Similarly, we can derive the conditions for the polarizaiton-switchable ECA between the $\mathbf{y}$ and $\mathbf{z}$ directions and between the $\mathbf{z}$ and $\mathbf{x}$ directions. Within effective mass approximation, we provide more general discussion on polarization-switchable ECA in Section A of the Supplementary Material (SM).
Excitingly, the polarization-switchable ECA implies the polarization-reversal induced conductivity change. For instance, the $\sigma_{x,x}(P,\mu)$ and $\sigma_{x,x}(-P,\mu)$ conductivity difference is given by
$\frac{\sigma_{x,x}(P,\mu)-\sigma_{x,x}(-P,\mu)}{\tau}\propto\frac{2\alpha_{1}P}{\sqrt{\alpha_{P}\beta_{P}\gamma_{P}}}\mu^{\frac{3}{2}},$ (6)
which suggests the asymmetry for resistances between $+P$ and $-P$ states, with this asymmetry depending on the magnitude of electric polarization.
Symmetry analysis. We now provide guidelines to screen nonpolar and nonmagnetic  [^4] As shown in Eqs. (3)–(6), neither magnetism nor spin-orbit interaction is essential for polarization-switchable ECA. Magnetism and spin-orbit interaction are thus neglected in our derivations. semiconductors enabling the polarization-switchable ECA. To this end, we focus on nonpolar crystallographic point groups, perform symmetry analysis, and derive symmetry-adapted energy dispersions for these groups. Depending on polarization $\mathbf{P}\equiv P_{x}\mathbf{x}+P_{y}\mathbf{y}+P_{z}\mathbf{z}$, the formalism of the energy dispersion is
$\epsilon(\mathbf{k},\mathbf{P})=\sum_{u,v=x,y,z}\tau_{uv}k_{u}k_{v}+\sum_{u,v,w=x,y,z}\lambda_{wuv}P_{w}k_{u}k_{v},$ (7)
where $\tau_{uv}$ and $\lambda_{wuv}$ coefficients characterize the bare effective masses and polarization-dressed effective masses of the carriers, respectively. Within effective mass approximation, Eq. (7) is derived around the Brillouin Zone center and up to first order in $P_{x}$, $P_{y}$ and $P_{z}$.
Table 1: The $\Gamma_{1}$ and $\Gamma_{4}$ irreducible representations for the $\bar{4}m2$ point group poi ; Koster et al. (1963); spa . In this table, $\mathfrak{1}$, $\bar{\mathfrak{4}}_{z}^{+}$, and $\bar{\mathfrak{4}}_{z}^{-}$ are identity operation, four-fold rotoinversion along $\mathbf{z}$ direction, and the inverse of $\bar{\mathfrak{4}}_{z}^{+}$, respectively; $\mathfrak{m}_{x}$ ($\mathfrak{m}_{y}$) is the mirror plane perpendicular to $\mathbf{x}$ ($\mathbf{y}$) direction, and $\mathfrak{2}_{xy}$ ($\mathfrak{2}_{x\bar{y}}$) is the two-fold rotation along $\mathbf{x}+\mathbf{y}$ ($\mathbf{x}-\mathbf{y}$) direction.
$\mathfrak{1}$ $\mathfrak{2}_{z}$ $\bar{\mathfrak{4}}_{z}^{+}$, $\bar{\mathfrak{4}}_{z}^{-}$ $\mathfrak{2}_{xy}$, $\mathfrak{2}_{x\bar{y}}$ $\mathfrak{m}_{x}$, $\mathfrak{m}_{y}$ Bases
$\Gamma_{1}$ $+1$ $+1$ $+1$ $+1$ $+1$ $k_{x}^{2}+k_{y}^{2}$, $k_{z}^{2}$
$\Gamma_{4}$ $+1$ $+1$ $-1$ $-1$ $+1$ $P_{z}$, $k_{x}^{2}-k_{y}^{2}$
In the following, we use $\bar{4}m2$ and $\bar{4}2m$ point groups to exemplify the derivations of the $P_{z}$-dependent energy dispersions. Our derivations are based on the group representation theory (see e.g., Refs. Hergert and Geilhufe (2018); El-Batanouny and Wooten (2008)). As shown in Table 1, the $k_{x}^{2}+k_{y}^{2}$ and $k_{z}^{2}$ bases belong to the identity representation $\Gamma_{1}$. This indicates that $(k_{x}^{2}+k_{y}^{2})$ and $k_{z}^{2}$ are invariant with respect to the $\bar{4}m2$ point group, yielding $\tau_{xx}(k_{x}^{2}+k_{y}^{2})$ and $\tau_{zz}k_{z}^{2}$ terms. The bases for the $\Gamma_{4}$ representation are $P_{z}$ and $k_{x}^{2}-k_{y}^{2}$, which are not invariant (e.g., $\bar{\mathfrak{4}}_{z}^{+}:P_{z}\rightarrow-P_{z},k_{x}^{2}-k_{y}^{2}\rightarrow-k_{x}^{2}+k_{y}^{2}$). Since $\Gamma_{4}\otimes\Gamma_{4}=\Gamma_{1}$, we can combine $P_{z}$ with $k_{x}^{2}-k_{y}^{2}$ to generate a new base for the $\Gamma_{1}$ representation [i.e., $\lambda_{zxx}P_{z}(k_{x}^{2}-k_{y}^{2})$]. Overall, the $P_{z}$-dependent energy dispersion associated with $\bar{4}m2$ is $\epsilon(\mathbf{k},P_{z})=(\tau_{xx}+\lambda_{zxx}P_{z})k_{x}^{2}+(\tau_{xx}-\lambda_{zxx}P_{z})k_{y}^{2}+\tau_{zz}k_{z}^{2}$. According to Eq. (5), this suggests the polarization-switchable ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions. Similarly, we can derive the $P_{z}$-dependent energy dispersion for the $\bar{4}2m$ point group as $\epsilon(\mathbf{k},P_{z})=\tau_{xx}(k_{x}^{2}+k_{y}^{2})+\tau_{zz}k_{z}^{2}+\lambda_{zxy}P_{z}k_{x}k_{y}$ [see “The $\bar{4}2m$ point group” in Section B of our SM]. Such an energy dispersion implies the polarization-switchable ECA between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions. To show this, we define an $x^{\prime}y^{\prime}z$ Cartesian coordinate system whose basis vectors are $\mathbf{x}^{\prime}=(\mathbf{x}+\mathbf{y})/\sqrt{2}$, $\mathbf{y}^{\prime}=(\mathbf{y}+\mathbf{y})/\sqrt{2}$, and $\mathbf{z}$ (see Fig. S1 of the SM). In the $x^{\prime}y^{\prime}z$ coordinate system, the electronic momentum is written as $\mathbf{k}=k_{x^{\prime}}\mathbf{x}^{\prime}+k_{y^{\prime}}\mathbf{y}^{\prime}+k_{z}\mathbf{z}$, and the energy dispersion associated with $\bar{4}2m$ becomes $\epsilon(\mathbf{k},P_{z})=(\tau_{xx}+\lambda_{zxy}P_{z}/2)k_{x^{\prime}}^{2}+(\tau_{xx}-\lambda_{zxy}P_{z}/2)k_{y^{\prime}}^{2}+\tau_{zz}k_{z}^{2}$. Therefore, the $\bar{4}2m$ point group hosts a polarization-switchable ECA between the $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$ (i.e., $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$) directions.
Table 2: Point groups enabling the polarization-switchable ECA, derived up to first order in electric polarization $P_{w}$ ($w=x,y,z$). The entries below each $P_{w}$ are endowed with $(\chi,\chi^{\prime})$ if ECA between the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions is enabled by gaining $P_{w}$ and is switchable by reversing $P_{w}$. Otherwise, the entries are endowed with $\cdots$”. Here, $\chi$ or $\chi^{\prime}$ is written as $u$, $uv$, or $\bar{u}v$ ($u,v$ being $x$, $y$, or $z$), where $uv$ and $\bar{u}v$ denote the $\mathbf{u}+\mathbf{v}$ and $-\mathbf{u}+\mathbf{v}$ directions, respectively.
Category A: rigorous cases
$P_{x}$ $P_{y}$ $P_{z}$
$\bar{4}$cdots $\cdots$ $(x,y)$; $(xy,\bar{x}y)$
$\bar{4}2m$cdots $\cdots$ $(xy,\bar{x}y)$
$\bar{4}m2$cdots $\cdots$ $(x,y)$
$312$ $(xy,\bar{x}y)$cdots $\cdots$
$321$cdots $(xy,\bar{x}y)$cdots
$\bar{6}m2$ $(xy,\bar{x}y)$cdots $\cdots$
$\bar{6}2m$cdots $(xy,\bar{x}y)$cdots
$23$ $(yz,\bar{y}z)$ $(zx,\bar{z}x)$ $(xy,\bar{x}y)$
$\bar{4}3m$ $(yz,\bar{y}z)$ $(zx,\bar{z}x)$ $(xy,\bar{x}y)$
Category B: approximate cases
$P_{x}$ $P_{y}$ $P_{z}$
$\bar{6}$ $(x,y)$; $(xy,\bar{x}y)$ $(x,y)$; $(xy,\bar{x}y)$cdots
$\bar{6}m2$cdots $(x,y)$cdots
$\bar{6}2m$ $(x,y)$cdots $\cdots$
$23$ $(\bar{y}z,x)$; $(x,yz)$ $(\bar{z}x,y)$; $(y,zx)$ $(\bar{x}y,z)$; $(z,xy)$
$\bar{4}3m$ $(\bar{y}z,x)$; $(x,yz)$ $(\bar{z}x,y)$; $(y,zx)$ $(\bar{x}y,z)$; $(z,xy)$
Our analyses for the entire nonpolar crystallographic point groups are shown in Section B and Section C of our SM. This helps to identify a variety of point groups that enable the polarization-switchable ECA, as summarized in Table 2. The cases listed in Table 2 can be classified into Category A and Category B. Belonging to Category A are cases where for ECA between the $\bm{\chi}$ and $\bm{\chi}^{\prime}$ directions (driven by $P_{w}$ polarization) the point group symmetry protects the $\sigma_{\chi,\chi}(P_{w},\mu)-\sigma_{\chi^{\prime},\chi^{\prime}}(P_{w},\mu)=-\sigma_{\chi,\chi}(-P_{w},\mu)+\sigma_{\chi^{\prime},\chi^{\prime}}(-P_{w},\mu)$ relationship. In such sense, the polarization-switchable feature for the ECA is rigorous, being robust against higher-order corrections  [^5] By higher-order corrections, we mean corrections that go beyond effective mass approximation or involve higher-order contributions from electric polarization. Such higher-order corrections should be compatible with the corresponding point group symmetry.. For instance, the $\bar{4}3m$ point group contains two-fold rotation symmetry operations along the $\mathbf{x}$ and $\mathbf{y}$ directions. These symmetry operations protect the polarization-switchable feature for the $P_{z}$-induced ECA between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions, because they provide a linkage between $\sigma_{xy,xy}(P_{z},\mu)-\sigma_{\bar{x}y,\bar{x}y}(P_{z},\mu)$ and $\sigma_{\bar{x}y,\bar{x}y}(-P_{z},\mu)-\sigma_{xy,xy}(-P_{z},\mu)$. The Category B contains cases for which the polarization-switchable ECA, derived within effective mass approximation (up to first-order in polarization), is not rigorously guaranteed by point group symmetry. As an example, the $\bar{6}2m$ point group hosts an ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions driven by $P_{x}$ polarization, while this point group has no symmetry operations linking $P_{x}$ with $-P_{x}$ (see Section C of the SM). This means that the $\sigma_{x,x}(P_{x},\mu)-\sigma_{y,y}(P_{x},\mu)=\sigma_{y,y}(-P_{x},\mu)-\sigma_{x,x}(-P_{x},\mu)$ relationship associated with the $\bar{6}2m$ point group may be broken when involving higher-order corrections such as $\zeta_{xyz}^{lmn}P_{x}^{l}k_{x}^{m}k_{y}^{n}$ ($l,m,n$ being natural numbers with $l\geq 2$ or $m+n>2$).
To end this section, we comment on the correlation between polarization-switchable ECA and polarization-reversal driven conductivity change. In Section A of the SM, we demonstrate that the $P_{w}$-switchable ECA between the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions is rooted in $(\eta_{\chi\chi}+\kappa_{w\chi\chi}P_{w})k_{\chi}^{2}$ or $(\eta_{\chi\chi}-\kappa_{w\chi\chi}P_{w})k_{\chi^{\prime}}^{2}$. According to Eqs. (3)–(6), $(\eta_{\chi\chi}+\kappa_{w\chi\chi}P_{w})k_{\chi}^{2}$ and $(\eta_{\chi\chi}-\kappa_{w\chi\chi}P_{w})k_{\chi^{\prime}}^{2}$ terms imply the polarization-reversal driven conductivity changes along the $\bm{\chi}$ and $\bm{\chi^{\prime}}$ directions, respectively. As for Table 2, we conclude that every $(\chi,\chi^{\prime})$ entry below $P_{w}$ indicates the conductivity changes resulting from the reversal of $P_{w}$, that is, $\sigma_{\chi,\chi}(P_{w})-\sigma_{\chi,\chi}(-P_{w})\neq 0$ and $\sigma_{\chi^{\prime},\chi^{\prime}}(P_{w})-\sigma_{\chi^{\prime},\chi^{\prime}}(-P_{w})\neq 0$. The only exceptions are associated with $23$ and $\bar{4}3m$ point groups, where $\sigma_{w,w}(P_{w})-\sigma_{w,w}(-P_{w})=0$ holds for $w$ being $x$, $y$, or $z$.
Figure 1: Schematizations of the crystal structures for $\bar{4}3m$ AlP (a) and $\bar{6}2m$ NaSrP (b). The $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ unit vectors in both panels are perpendicular to each other. In panel (a), the $\mathbf{x^{\prime}}$ and $\mathbf{y^{\prime}}$ unit vectors are defined as $\mathbf{x^{\prime}}=(\mathbf{x}+\mathbf{y})/\sqrt{2}$ and $\mathbf{y^{\prime}}=(-\mathbf{x}+\mathbf{y})/\sqrt{2}$. The boxes enclosed by cyan solid lines (i.e., $xyz$ coordinate system) and black dashed lines (i.e., $x^{\prime}y^{\prime}z$ coordinate system) denote the unit cell of AlP and the super cell used in our simulations, respectively.
Semiconductors with polarization-switchable ECA. We move on to verify our theory and identify semiconductors with polarization-switchable ECA, by electrical conductivity calculations based on first principles. As shown in Table 2, a variety of point groups host the ECA driven by polarization and rigorously switched by flipping the polarization. Of particular interest is the $\bar{4}3m$ point group, where a majority of technologically important semiconductors (i.e., zinc-blende type) belong to this point group Hinuma et al. (2014); Grüneis et al. (2014). In the absence of electric polarization, the electrical conductivities associated with the $\bar{4}3m$ point group is isotropic among the $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ directions. When gaining a polarization along $\mathbf{z}$, the rigorously polarization-switchable ECA is enabled between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions (see Table 2). To verify this, we take the zinc-blende type AlP semiconductor, with a band gap of $\sim$2.4 eV Richman (1968), as our platform. In Fig. 1(a), the box enclosed by cyan solid lines schematizes the unit cell of AlP (i.e., $xyz$ coordinate system), while the larger super cell (i.e., $x^{\prime}y^{\prime}z$ coordinate system) is the one employed in our numerical simulations. To polarize the AlP semiconductor, we apply external electric fields of ±$\pm 2$ and ±$\pm 4$ MV/cm along ±$\pm\mathbf{z}$ direction to such a material. This creates four different $P_{z}$ polarization states in AlP, and yields finite $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})/\tau$ conductivity differences being $P_{z}$ dependent [see Fig. 2(a)]. According to Fig. 1(a), such conductivity differences characterize the ECA between the $\mathbf{x}+\mathbf{y}$ and $-\mathbf{x}+\mathbf{y}$ directions. Remarkably, the conductivity differences associated with electric fields of $+4$ and $-4$ MV/cm ($+2$ and $-2$ MV/cm) are rather symmetric with respect to the horizontal axis. This confirms our theory that $P_{z}$ drives strictly polarization-switchable ECA in materials with $\bar{4}3m$ point group.
Table 2 also contains several cases associated with approximately polarization-switchable ECA. A representative case is the $\bar{6}2m$ point group which hosts ECA between the $\mathbf{x}$ and $\mathbf{y}$ directions — being driven by $P_{x}$ electric polarization but not rigorously switchable by reversing $P_{x}$. We confirm this by computing the electrical conductivities for the NaSrP semiconductor with $\bar{6}2m$ symmetry Dong and DiSalvo (2007). The crystal structure of NaSrP is schematized in Fig. 1(b). As shown in Fig. S2 of the SM, both the valence band maximum (VBM) and conduction band minimum (CBM) for NaSrP locate at the Brillouin Zone center. This implies that the $P_{x}$-dependent energy dispersion $\epsilon(\mathbf{k},P_{x})=\tau_{xx}(k_{x}^{2}+k_{y}^{2})+\tau_{zz}k_{z}^{2}+\lambda_{xxx}P_{x}(k_{x}^{2}-k_{y}^{2})$ [see Eq. S(49) of the SM] should be approximately valid for chemical potential $\mu$ being close to VBM or CBM. To create $P_{x}$ in NaSrP, we polarize it by applying external electric fields of ±$\pm 4$ MV/cm along ±$\pm\mathbf{x}$ direction. The resulting $(\sigma_{x,x}-\sigma_{y,y})/\tau$ conductivity differences, associated with electric fields of ±$\pm 4$ MV/cm, are basically symmetric with respect to the horizontal axis [see Fig. 2(b)]; The asymmetric feature becomes significant for $\mu$ being larger than 2.8 eV. This validates that the $\bar{6}2m$ point group enables approximately polarization-switchable ECA.
We further numerically check the polarization-reversal induced conductivity changes in AlP and NaSrP. As shown in Fig. S3 of the SM, the $+2$ MV/cm and $-2$ MV/cm electric fields yield diverse $\sigma_{x^{\prime},x^{\prime}}$ conductivities in AlP, which indicates conductivity changes induced by polarization reversal (associated with ±$\pm 2$ MV/cm electric fields). Consistent with Eq. (6), the conductivity changes associated with ±$\pm 4$ MV/cm electric fields are almost doubled compared with those associated with ±$\pm 2$ MV/cm electric fields. Similarly, the polarization-reversal induced conductivity changes are also predicted to occur in NaSrP (see Fig. S4 of the SM).
Figure 2: The ECA in AlP and NaSrP as functions of chemical potential $\mu$, where $\mu=0$ is referred to as the corresponding valence band maximum. Panel (a): $(\sigma_{x^{\prime},x^{\prime}}-\sigma_{y^{\prime},y^{\prime}})/\tau$ in AlP driven by electric fields along ±$\pm\mathbf{z}$ direction. Panel (b): $(\sigma_{x,x}-\sigma_{y,y})/\tau$ in NaSrP driven by electric fields along ±$\pm\mathbf{x}$ direction. In panels (a) and (b), the unit for the vertical axes is $10^{19}/(\Omega\cdot m\cdot s)$.
Summary and outlook. Within effective mass approximation, we have developed a theory on polarization-switchable ECA in nonpolar semiconductors, where the ECA exhibits a first-order response to polarization and is reversible by flipping the polarization. We have shown that the polarization-switchable ECA is accompanied with conductivity changes driven by the reversal of the electric polarization. Hosted by a variety of point groups (summarized in Table 2), our discussed polarization-switchable ECA phenomena can be classified into two categories. Associated with the first category are cases for which the ECA is rigorously switchable by reversing the polarization. Such a polarization-switchable feature is protected by point group symmetry and is robust against higher-order corrections (e.g., beyond effective mass approximation). In contrast, the polarization-switchable feature for ECA belonging to the second category — valid in the regime of effective mass approximation and first-order response to polarization — may be broken when involving higher-order corrections. Our theory is further numerically validated by electrical conductivity calculations based on first principles. This yields the predictions of AlP and NaSrP as two representative semiconductors with polarization-switchable ECA and polarization-reversal driven conductivity change. We expect that our theory opens a route toward the design of next-generation electronic devices based on our aforementioned anisotropic charge transport and related phenomena.
Acknowledgements. – We acknowledge the support from the National Natural Science Foundation of China (Grants Nos. 12274174, T2225013, 52288102, 52090024, and 12034009). L.B. thanks the Vannevar Bush Faculty Fellowship (VBFF) grant No. N00014-20-1-2834 from the Department of Defense and award No. DMR-1906383 from the National Science Foundation AMASE-i Program (MonArk NSF Quantum Foundry). H.J.Z. thanks the supports from “Xiaomi YoungScholar” Project, Arkansas High Performance Computing Center, and high-performance computing center of Jilin University.